OpenStudy (anonymous):

how would you solve log base 2 of (X^2- 6x)= 4

OpenStudy (turingtest):

\[\log_2(x^2-6x)=4\]Raise 2 to the power of both sides and you will have a very solvable quadratic equation.

OpenStudy (anonymous):

thanks btw can you stay on so i can ask u more questions if i need help

OpenStudy (anonymous):

x=-1?

OpenStudy (turingtest):

sure I'll be here a minute, but I have other people to help too.

OpenStudy (anonymous):

alright how would you solve (4)^ -3/4x=1/2

OpenStudy (anonymous):